# An Introduction to Single-Antenna Radio Astronomical Polarimetry

### v2

@@ 4.3 Mueller Matrices
$M_{\mu}^{\nu}=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\left({\bf J}\mbox{$\sigma$}_{\nu}\,{\bf J}^{\dagger}\right)}.$ (41)
Proof 4.5.
$S_{\mu}^{\prime}$ $={\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\mbox{$\rho$}^{\prime}}$ Eqn. (16) $={\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\left({\bf J}\mbox{$\rho$}\,{\bf J}^{\dagger}\right)}$ Eqn. (37) $={\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\left({\bf J}\left[S_{\nu}\,\mbox{$\sigma$}_{\nu}/2\right]{\bf J}^{\dagger}\right)}$ Eqn. (14) $=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\left({\bf J}\mbox{$\sigma$}_{\nu}\,{\bf J}^{\dagger}\right)}S_{\nu}$ $S_{\nu}$ is a scalar $=M_{\mu}^{\nu}S_{\nu}$ Eqn. (41)
@@
The signal path described by van Straten et al. (2010) extends the Hamaker et al. (1996) model with an impure correction for the complex phase convention, which is described in more detail in Section 5.1. Therefore, the van Straten et al. (2010) calibration model is best summarized using the equivalent transformation of the Stokes parameters by a Mueller matrix $\underline{\underline{{\bf M}}}$,
@@
From right to left, $\underline{\underline{\mbox{$S$}}}$ is the column vector with elements equal to the 4 Stokes parameters intrinsic to the astronomical source; $\underline{\underline{{\bf F}}}$ through $\underline{\underline{{\bf G}}}$ are the Mueller matrices derived from the Jones matrices defined in equation (36) and discussed in Section 5; $\underline{\underline{{\bf\Phi}}}$ is the phase convention correction (see Section 5.1); and $\underline{\underline{\mbox{$S$}}}^{\prime}$ is the column vector of observed Stokes parameters.
4.4 Polarization Transfer Tensors
@@
$\mbox{$\rho$}^{\prime}=\mbox{$\mathcal{U}$}:\mbox{$\rho$}.$ (43)
@@
$\left({\bf A}\otimes{\bf B}\right)\mbox{$:$}\,{\bf C}$ $=$ ${\bf A}\left({\bf B}\,\mbox{$:$}\,{\bf C}\right)$ (46)
$\left({\bf A}\,\tilde{\otimes}\,{\bf B}\right)\mbox{$:$}\,{\bf C}$ $=$ ${\bf A}{\bf C}{\bf B}$ (47)
@@
Proof 4.6.
$\mbox{$\rho$}^{\prime}$ $=\mbox{$\mathcal{U}$}:\mbox{$\rho$}$ Eqn. (43) $=\left({\bf J}\,\tilde{\otimes}\,{\bf J}^{\dagger}\right):\mbox{$\rho$}$ Eqn. (48) $={\bf J}\,\mbox{$\rho$}\,{\bf J}^{\dagger}.$ Eqn. (47)
@@
$M_{\mu}^{\nu}=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\mbox{$\mathcal{U}$}}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}}.$ (49)
@@
Proof 4.7.
$\frac{1}{2}$ ${\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\mbox{$\mathcal{U}$}}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}}$ Eqn. (49) $=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\left(\frac{1}{2}M_{\kappa}^{\lambda}{\mbox{$\sigma$}_{\kappa}}\otimes{\mbox{$\sigma$}_{\lambda}}\right)}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}}$ Eqn. (50) $=\frac{1}{4}M_{\kappa}^{\lambda}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\left({\mbox{$\sigma$}_{\kappa}}\otimes{\mbox{$\sigma$}_{\lambda}}\right)}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}}$ $=\frac{1}{4}M_{\kappa}^{\lambda}\left({\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\kappa}}\right)\left({\mbox{$\sigma$}_{\lambda}}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}\right)$ Eqn. (46) $=M_{\kappa}^{\lambda}\delta_{\mu\kappa}\delta_{\nu\lambda}$ Eqn. (C4) $=M_{\mu}^{\nu}$
@@
4.5 Comparison with Radio Interferometry
@@
$\mbox{$V$}_{AB}={\mbox{$\mathcal{U}$}_{AB}}\,\mbox{$:$}\,{\mbox{$\rho$}},$ (53)
@@ 4.6 Polar Decomposition
Polar decomposition into Hermitian and unitary matrices provides a practical framework for classifying and conceptualizing distinct polarization-altering effects, as discussed in more detail in Section 4.7. Isolating the determinant of the Jones matrix also has practical benefit. In single-antenna polarimetry, the coherency matrix is insensitive to the phase of $J$ (Proof 4.2); therefore, without any loss of generality, the absolute phase is set to zero and $J$ is replaced by the real-valued absolute gain, $G=|J|$, thereby eliminating a degenerate dof.
@@
4.7 Axis-Angle Representation
@@ 5.2 Receiver Gains
a boost component ${\bf B}_{g}$, and a rotation component ${\bf R}_{g}$. As noted in Section 4.6, the absolute phase is chosen such that $G$ is real-valued. The following sub-sections focus on the axis-angle representations of ${\bf B}_{g}$ and ${\bf R}_{g}$, where it is shown that ${\bf B}_{g}$ characterizes the ratio of the gain amplitudes, or differential gain, and ${\bf R}_{g}$ describes the differential phase.
5.2.1 Differential Gain
@@
To arrive at the axis-angle representation of ${\bf B}_{g}$, note that its off-diagonal elements are zero and only $\mbox{$\sigma$}_{0}$ and $\mbox{$\sigma$}_{1}$ contribute to the diagonal elements of the sum in equation (59). Therefore, $\hat{\underline{\mbox{$m$}}}=(1,0,0)$ and
@@
5.2.2 Differential Phase
The axis-angle representation of ${\bf R}_{g}$ is obtained by rearranging the polar decomposition of G to arrive at ${\bf R}_{g}=G^{-1}{\bf B}_{g}^{-1}{\bf G}$. Both G and ${\bf B}_{g}$ (and their inverses) are diagonal and, as a product of diagonal matrices, ${\bf R}_{g}$ is also diagonal. Only $\mbox{$\sigma$}_{0}$ and $\mbox{$\sigma$}_{1}$ contribute to the diagonal elements of the sum in equation (60); therefore, $\hat{\underline{\mbox{$n$}}}=(1,0,0)$ and
@@ 5.2.2 Differential Phase
Differential phase describes any transformation that subjects opposite polarizations to different delays; it is also known as retardance in optics and cross-hand phase or delay in interferometry. It arises when the opposite polarizations propagate along signal paths of different lengths and/or with different speeds. For a given delay $\Delta t$ between $e_{0}(t)$ and $e_{1}(t)$, the corresponding differential phase $\phi=\nu\Delta t$ varies linearly with radio frequency. Nonlinear spectral variations arise from dispersive effects in the electronics of the receiver, down-conversion system, cables, and other components used to connect observatory instrumentation. Section 5.6 discusses the differential phase that arises during propagation through the magnetized plasma in the Earth’s ionosphere and the interstellar medium.
5.3 Deviations from an Ideal Feed
@@ 5.3.1 Non-orthogonal Component
Let ${\bf B}_{d}^{2}={\bf B}_{\hat{\underline{\mbox{\scriptsize$m$}}}}(2\beta)$ and note that the elements on the diagonal of equation (72) are equal to each other; therefore, there can be no contribution from $\mbox{$\sigma$}_{1}$. Only $\mbox{$\sigma$}_{2}$ and $\mbox{$\sigma$}_{3}$ contribute to the off-diagonal elements of the sum in equation (59); i.e., $\hat{\underline{\mbox{$m$}}}=(0,m_{2},m_{3})$, and
@@
As in Section 4.1 of van Straten (2004), the 2 dof in the non-orthogonal component of D may be parameterized by $\underline{\mbox{$b$}}=(0,b_{2},b_{3})=\hat{\underline{\mbox{$m$}}}\sinh\beta$, such that $\sinh^{2}\beta=\underline{\mbox{$b$}}\cdot\underline{\mbox{$b$}},$ $\cosh^{2}\beta=1+\underline{\mbox{$b$}}\cdot\underline{\mbox{$b$}},$ and
@@ 5.3.2 Orthonormal Component
$\underline{\mbox{$S$}}=I\left(\cos 2\chi\cos 2\psi,\;\cos 2\chi\sin 2\psi,\;\sin 2\chi\right).$ (80)
@@
and the polarization vector $\underline{\mbox{$S$}}^{\prime}=I(1,0,0)$.
@@
5.4 Nominal feed configuration
@@
Feed Basis LIN or CIRC $\underline{\mbox{$S$}}$=(Q,U,V) $\underline{\mbox{$S$}}$=(V,Q,U)
@@
5.4.1 Feed Basis
Typically, the two receptors in a receiver are either linearly or circularly polarized, and Section 2 describes a Cartesian coordinate system defined by linearly polarized receptors. As shown in Appendix D.4, the transformation from this Cartesian basis to one defined by a pair of circularly-polarized receptors is described by
@@ 5.4.1 Feed Basis
In the basis defined by ${\bf R}_{b}$, the Stokes polarization vector $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})=(V,Q,U)$ and the effects of all subsequent transformations (from X to $\underline{\underline{{\bf\Phi}}}$) differ from their impact in the original Cartesian basis.
@@
In the circular basis, the differential phase $\phi$ describes rotation about the Stokes V axis, which is equivalent to a physical rotation about the line of sight (Proof 5.2). Later sections describe the impacts of the circular basis on the feed hand ${\bf R}_{h}$ (Section 5.4.2), axis of symmetry ${\bf R}_{\hat{\underline{\mbox{\scriptsize$z$}}}}(\Theta)$ (Section 5.4.3), and phase convention $\underline{\underline{{\bf\Phi}}}$ (Section 6.1) transformations.
5.4.2 Feed Hand / Basis Reflection
@@
5.4.3 Axis of Symmetry
The axis of symmetry in the $x$-$y$ plane depicted in Figure 1 is defined as the position angle of a linearly-polarized wave that produces equal responses in each ideal receptor. In a basis defined by linearly-polarized receptors, a linearly-polarized wave that oscillates along $x=y$ (positive Stokes $U$) will produce equal and in-phase responses in each receptor; therefore, the symmetry angle has a nominal value of 45. In a basis defined by circularly-polarized receptors, a linearly-polarized wave that oscillates along the $x$-axis will produce equal and in-phase responses in each receptor, and the symmetry angle has a nominal value of 0.
The symmetry axis of a receiver on steerable mount describes its rotation about the line of sight with respect to a reference point on the antenna structure. For a fixed dipole on the ground, the symmetry axis describes its rotation about the zenith. Once established, the symmetry angle is typically treated as a constant. Any unintended rotation (either geometric or apparent; e.g., owing to cross-coupling of the receptors) must be determined through calibration of the orthonormal component of deviation, which includes an orientation angle $\psi$ (eqn. 82). Other known rotations are included in the projection transformation described in Section 5.5.
Figure 5: In the circular basis, the direction of the Stokes polarization vector, $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})=(V,Q,U)$, is defined by the polar angles, $\xi$ and $\zeta$. These angles together define a right triangle on the surface of the Poincaré sphere with hypotenuse equal to the colatitude, $\theta=\pi/2-2\chi$, measured with respect to the pole defined by positive $S_{1}$ (Stokes V). Also depicted is the longitude $2\psi$, measured with respect to positive $S_{2}$ (Stokes Q) toward positive $S_{3}$ (Stokes U).
5.5 Projection onto the Celestial Sphere

### v4

@@ IV.3 Mueller Matrices
$M_{\mu}^{\nu}=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{\left({\bf J}\mbox{$\sigma$}_{\nu}\,{\bf J}^{\dagger}\right)}.$ (41)
Proof IV.5.
$S_{\mu}^{\prime}$ $={\mbox{$\sigma$}_{\mu}}\mathbin{:}{\mbox{$\rho$}^{\prime}}$ Eqn. (16) $={\mbox{$\sigma$}_{\mu}}\mathbin{:}{\left({\bf J}\mbox{$\rho$}\,{\bf J}^{\dagger}\right)}$ Eqn. (37) $={\mbox{$\sigma$}_{\mu}}\mathbin{:}{\left({\bf J}\left[S_{\nu}\,\mbox{$\sigma$}_{\nu}/2\right]{\bf J}^{\dagger}\right)}$ Eqn. (14) $=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{\left({\bf J}\mbox{$\sigma$}_{\nu}\,{\bf J}^{\dagger}\right)}S_{\nu}$ $S_{\nu}$ is a scalar $=M_{\mu}^{\nu}S_{\nu}$ Eqn. (41)
@@
The signal path described by van Straten et al. (2010) extends the Hamaker et al. (1996) model with an impure correction for the complex phase convention, which is described in more detail in Section V.1. Therefore, the van Straten et al. (2010) calibration model is best summarized using the equivalent transformation of the Stokes parameters by a Mueller matrix $\underline{\underline{{\bf M}}}$,
@@
From right to left, $\underline{\underline{\mbox{$S$}}}$ is the column vector with elements equal to the 4 Stokes parameters intrinsic to the astronomical source; $\underline{\underline{{\bf F}}}$ through $\underline{\underline{{\bf G}}}$ are the Mueller matrices derived from the Jones matrices defined in Equation (36) and discussed in Section V; $\underline{\underline{{\bf\Phi}}}$ is the phase convention correction (see Section V.1); and $\underline{\underline{\mbox{$S$}}}^{\prime}$ is the column vector of observed Stokes parameters.
IV.4 Polarization Transfer Tensors
@@
$\mbox{$\rho$}^{\prime}=\mbox{$\mathcal{U}$}\mathbin{:}\mbox{$\rho$}.$ (43)
@@
$\left({\bf A}\otimes{\bf B}\right)\mathbin{:}{\bf C}$ $=$ ${\bf A}\left({\bf B}\mathbin{:}{\bf C}\right)$ (46)
$\left({\bf A}\,\tilde{\otimes}\,{\bf B}\right)\mathbin{:}{\bf C}$ $=$ ${\bf A}{\bf C}{\bf B}$ (47)
@@
Proof IV.6.
$\mbox{$\rho$}^{\prime}$ $=\mbox{$\mathcal{U}$}\mathbin{:}\mbox{$\rho$}$ Eqn. (43) $=\left({\bf J}\,\tilde{\otimes}\,{\bf J}^{\dagger}\right)\mathbin{:}\mbox{$\rho$}$ Eqn. (48) $={\bf J}\,\mbox{$\rho$}\,{\bf J}^{\dagger}.$ Eqn. (47)
@@
$M_{\mu}^{\nu}=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\mbox{$\mathcal{U}$}}\mathbin{:}{\mbox{$\sigma$}_{\nu}}}.$ (49)
@@
Proof IV.7.
$\frac{1}{2}$ ${\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\mbox{$\mathcal{U}$}}\mathbin{:}{\mbox{$\sigma$}_{\nu}}}$ Eqn. (49) $=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\left(\frac{1}{2}M_{\kappa}^{\lambda}{\mbox{$\sigma$}_{\kappa}}\otimes{\mbox{$\sigma$}_{\lambda}}\right)}\mathbin{:}{\mbox{$\sigma$}_{\nu}}}$ Eqn. (50) $=\frac{1}{4}M_{\kappa}^{\lambda}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\left({\mbox{$\sigma$}_{\kappa}}\otimes{\mbox{$\sigma$}_{\lambda}}\right)}\mathbin{:}{\mbox{$\sigma$}_{\nu}}}$ $=\frac{1}{4}M_{\kappa}^{\lambda}\left({\mbox{$\sigma$}_{\mu}}\mathbin{:}{\mbox{$\sigma$}_{\kappa}}\right)\left({\mbox{$\sigma$}_{\lambda}}\mathbin{:}{\mbox{$\sigma$}_{\nu}}\right)$ Eqn. (46) $=M_{\kappa}^{\lambda}\delta_{\mu\kappa}\delta_{\nu\lambda}$ Eqn. (C4) $=M_{\mu}^{\nu}$
@@
IV.5 Comparison with Radio Interferometry
@@
$\mbox{$V$}_{AB}={\mbox{$\mathcal{U}$}_{AB}}\mathbin{:}{\mbox{$\rho$}},$ (53)
@@ IV.6 Polar Decomposition
Polar decomposition into Hermitian and unitary matrices provides a practical framework for classifying and conceptualizing distinct polarization-altering effects, as discussed in more detail in Section IV.7. Isolating the determinant of the Jones matrix also has practical benefit. In single-antenna polarimetry, the coherency matrix is insensitive to the phase of $J$ (Proof IV.2); therefore, without any loss of generality, the absolute phase is set to zero and $J$ is replaced by the real-valued absolute gain, $G=|J|$, thereby eliminating a degenerate dof.
@@
IV.7 Axis-Angle Representation
@@ V.2 Receiver Gains
a boost component ${\bf B}_{g}$, and a rotation component ${\bf R}_{g}$. As noted in Section IV.6, the absolute phase is chosen such that $G$ is real-valued. The following sub-sections focus on the axis-angle representations of ${\bf B}_{g}$ and ${\bf R}_{g}$, where it is shown that ${\bf B}_{g}$ characterizes the ratio of the gain amplitudes, or differential gain, and ${\bf R}_{g}$ describes the differential phase.
V.2.1 Differential Gain
@@
To arrive at the axis-angle representation of ${\bf B}_{g}$, note that its off-diagonal elements are zero and only $\mbox{$\sigma$}_{0}$ and $\mbox{$\sigma$}_{1}$ contribute to the diagonal elements of the sum in Equation (59). Therefore, $\hat{\underline{\mbox{$m$}}}=(1,0,0)^{T}$ and
@@
V.2.2 Differential Phase
The axis-angle representation of ${\bf R}_{g}$ is obtained by rearranging the polar decomposition of G to arrive at ${\bf R}_{g}=G^{-1}{\bf B}_{g}^{-1}{\bf G}$. Both G and ${\bf B}_{g}$ (and their inverses) are diagonal and, as a product of diagonal matrices, ${\bf R}_{g}$ is also diagonal. Only $\mbox{$\sigma$}_{0}$ and $\mbox{$\sigma$}_{1}$ contribute to the diagonal elements of the sum in Equation (60); therefore, $\hat{\underline{\mbox{$n$}}}=(1,0,0)^{T}$ and
@@ V.2.2 Differential Phase
Differential phase describes any transformation that subjects opposite polarizations to different delays; it is also known as retardance in optics and cross-hand phase or delay in interferometry. It arises when the opposite polarizations propagate along signal paths of different lengths and/or with different speeds. For a given delay $\Delta t$ between $e_{0}(t)$ and $e_{1}(t)$, the corresponding differential phase $\phi=\nu\Delta t$ varies linearly with radio frequency. Nonlinear spectral variations arise from dispersive effects in the electronics of the receiver, down-conversion system, cables, and other components used to connect observatory instrumentation. Section V.6 discusses the differential phase that arises during propagation through the magnetized plasma in the Earth’s ionosphere and the interstellar medium.
V.3 Deviations from an Ideal Feed
@@ V.3.1 Non-orthogonal Component
Let ${\bf B}_{d}^{2}={\bf B}_{\hat{\underline{\mbox{\scriptsize$m$}}}}(2\beta)$ and note that the elements on the diagonal of Equation (72) are equal to each other; therefore, there can be no contribution from $\mbox{$\sigma$}_{1}$. Only $\mbox{$\sigma$}_{2}$ and $\mbox{$\sigma$}_{3}$ contribute to the off-diagonal elements of the sum in Equation (59); i.e., $\hat{\underline{\mbox{$m$}}}=(0,m_{2},m_{3})^{T}$, and
@@
As in Section 4.1 of van Straten (2004), the 2 dof in the non-orthogonal component of D may be parameterized by $\underline{\mbox{$b$}}=(0,b_{2},b_{3})^{T}=\hat{\underline{\mbox{$m$}}}\sinh\beta$, such that $\sinh^{2}\beta=\underline{\mbox{$b$}}\cdot\underline{\mbox{$b$}},$ $\cosh^{2}\beta=1+\underline{\mbox{$b$}}\cdot\underline{\mbox{$b$}},$ and
@@ V.3.2 Orthonormal Component
$\underline{\mbox{$S$}}=I\left(\cos 2\chi\cos 2\psi,\;\cos 2\chi\sin 2\psi,\;\sin 2\chi\right)^{T}.$ (80)
@@
and the polarization vector $\underline{\mbox{$S$}}^{\prime}=I(1,0,0)^{T}$.
@@
V.4 Nominal Feed Configuration
@@
Feed Basis LIN or CIRC $\underline{\mbox{$S$}}=(Q,U,V)^{T}$ $\underline{\mbox{$S$}}=(V,Q,U)^{T}$
@@
V.4.1 Feed Basis
Typically, the two receptors in a receiver are either linearly or circularly polarized, and Section II describes a Cartesian coordinate system defined by linearly polarized receptors. As shown in Appendix D.4, the transformation from this Cartesian basis to one defined by a pair of circularly-polarized receptors is described by
@@ V.4.1 Feed Basis
In the basis defined by ${\bf R}_{b}$, the Stokes polarization vector $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})^{T}=(V,Q,U)^{T}$ and the effects of all subsequent transformations (from X to $\underline{\underline{{\bf\Phi}}}$) differ from their impact in the original Cartesian basis.
@@
In the circular basis, the differential phase $\phi$ describes rotation about the Stokes V axis, which is equivalent to a physical rotation about the line of sight (Proof V.2). Later sections describe the impacts of the circular basis on the feed hand ${\bf R}_{h}$ (Section V.4.2), axis of symmetry ${\bf R}_{\hat{\underline{\mbox{\scriptsize$z$}}}}(\Theta)$ (Section V.4.3), and phase convention $\underline{\underline{{\bf\Phi}}}$ (Section VI.1) transformations.
V.4.2 Feed Hand / Basis Reflection
@@
V.4.3 Axis of Symmetry
The axis of symmetry in the $x$-$y$ plane depicted in Figure 1 is defined as the position angle of a linearly-polarized wave that produces equal responses in each ideal receptor. In a basis defined by linearly-polarized receptors, a linearly-polarized wave that oscillates along $x=y$ (positive Stokes $U$) will produce equal and in-phase responses in each receptor; therefore, the symmetry angle has a nominal value of 45. In a basis defined by circularly-polarized receptors, a linearly-polarized wave that oscillates along the $x$ axis will produce equal and in-phase responses in each receptor, and the symmetry angle has a nominal value of 0.
The symmetry axis of a receiver on steerable mount describes its rotation about the line of sight with respect to a reference point on the antenna structure. For a fixed dipole on the ground, the symmetry axis describes its rotation about the zenith. Once established, the symmetry angle is typically treated as a constant. Any unintended rotation (either geometric or apparent; e.g., owing to cross-coupling of the receptors) must be determined through calibration of the orthonormal component of deviation, which includes an orientation angle $\psi$ (Eqn. 82). Other known rotations are included in the projection transformation described in Section V.5.
Figure 5: In the circular basis, the direction of the polarization vector, $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})^{T}=(V,Q,U)^{T}$, is defined by the polar angles, $\xi$ and $\zeta$. These angles together define a right triangle on the surface of the Poincaré sphere with hypotenuse equal to the colatitude, $\theta=\pi/2-2\chi$, measured with respect to the pole defined by positive $S_{1}$ (Stokes V). Also depicted is the longitude $2\psi$, measured with respect to positive $S_{2}$ (Stokes Q) toward positive $S_{3}$ (Stokes U).
V.5 Projection onto the Celestial Sphere